OpenStudy (luigi0210):

A lake is stocked with 500 fish and their population \(t\) months after the fish were introduced is \(\large p(t)=\frac{10,000}{1+19e^{-0.2t}}\). Observe that \(p(0)=500\). At what rate is the fish population changing, one month after the lake is stocked?

OpenStudy (luigi0210):

@agent0smith @whpalmer4 @jigglypuff314

mathslover (mathslover):

You need to find \(\cfrac{dP(t)}{dt} \) here. Right?

OpenStudy (luigi0210):

Uh, I think so.

mathslover (mathslover):

We will solve this by using chain rule.

mathslover (mathslover):

Let us take, the denominator as "x" So, differentiation of \(\cfrac{10^4}{x} = ? \)

OpenStudy (luigi0210):

\[\LARGE -\frac{10^4}{x^2}\]

mathslover (mathslover):

Right! And now we will differentiate the rest part. Differentiation of \(1+ 19e^{-0.2t}\) = ?

OpenStudy (luigi0210):

\[\LARGE (-0.2)(19e^{-0.2t})\]

mathslover (mathslover):

Right so you have : \(\cfrac{-10^4}{(1+19e^{-0.2t})^2} \times (-0.2) \times (19e^{-0.2t})\)

OpenStudy (luigi0210):

What do we do from here? o.O

mathslover (mathslover):

Hmm, I think just simplifying this further is enough for us? \(\cfrac{3.8 \times 10^4 \times 19e^{-0.2t} }{ (1+19e^{-0.2t})^2}\)

mathslover (mathslover):

\(\cfrac{3.8 \times 10^4 \times e^{-0.2t} }{ (1+19e^{-0.2t})^2}\) - Sorry, I didn't remove 19 from there.